jennifer is one-fourth her father's age. In twelve years, Jennifer will be the same age as her father was 15 years ago. find the present ages of jennifer and her father.
Answer by mathslover(157)

Let the present age of Jennifer's father be x
therefore Jennifer's present age = x/4
In twelve years Jennifer's age = x/4 + 12
15 years ago Jeniffer's fathers age was
x-15
Since both these ages are same
=> x/4 + 12 = x-15
Transpose the similar terms together
x - x/4 = 15 + 12
3x/4 = 27
Multiplying by 4/3 on both sides
x= 27* 4/3 =36
Jeniffer's father's age =36
Jeniffer's age = 36/4 = 9